In the Bernoulli's equation prove that 1/2 V^2 has same dimension as pressure.
1
Expert's answer
2019-05-23T09:43:53-0400

The dimensions are


"[\rho] = \frac{M}{L^3}"

"[v] = \frac{L}{T}"

hence


"\bigg[\frac{\rho v^2}{2}\bigg] = \frac{M}{L^3} \frac{L^2}{T^2} = M L^{-1} T^{-2}"

On the other hand


"[p] = \bigg[\frac{F}{S}\bigg] = \frac{M L}{T^2} \frac{1}{L^2} = M L^{-1} T^{-2}"

The dimensions are the same.